ltem 1 I ReviewI Constants | Periodic Table Part A A baseball pitcher can throw a ball with a speed of 70 m/s. He is in the bA baseball pitcher can throw a ball with a speed of 70 m/s . He is in the back of a pickup truck that is driving away from you. He throws the ball in your direction, and it floats toward you at a lazy 10 m/s
Answer #1

The ball has two velocities (i) given by the truck in whichthe pitcher is travelling., vt

+(ii) the velocity given by the pitvher, vp = 70 m/s

Both are along the same line but in opposite direction.

Resultant Velocity of the ball,

                  Vb = Vp - Vt

Hence,       Vt =Vp - Vb = 70 - 10 = 60 m/s
Answer #2

ANSWER :


Truck carrying pitcher is moving away from me.


Ball is thrown by the pitcher from the moving truck towards me at a speed of 70 m/s.


Ball in reality travels at a speed of 10 m/s towards me.


This is because of movement of truck and ball are in opposite directions.


Let speed of truck be V m/s. So, ball acquires the same speed and direction.

If ball is thrown in my direction at 70 m/s and it is moving away from me at V m/s, then, 

ball ’s real speed towards me = 70 - V   m/s


But it is noted as 10 m/s.


So,


70 - V = 10

=> V = 70 - 10 = 60 m/s.


So, speed of truck = 60 m/s (ANSWER).

answered by: Tulsiram Garg